OpenStudy (anonymous):

is 7x-1/6x a polynomial, how many terms and what is the degree

OpenStudy (agent0smith):

Hard to interpret what it is... use parentheses. If it's \[\Large \frac{ 7x-1 }{6x }\]or \[\Large7x- \frac{ 1 }{6x }\] Then it is not a polynomial. A polynomial can't have a variable in the denominator. This is a polynomial:\[\Large7x- \frac{ 1 }{6 }x\]

OpenStudy (anonymous):

IS IT A POLYNOMIAL

OpenStudy (anonymous):

HOW MANY TERMS

OpenStudy (anonymous):

ARE YOU STILL THERE

OpenStudy (mathstudent55):

Is it \( 7x - \dfrac{1}{6} x \) or \( 7x - \dfrac{1}{6x} \) The first one is a polynomial; the second one isn't.

OpenStudy (agent0smith):

@ONIONLDT I tried to answer your question. We can't unless you clarify it.